### v1

@@ 5.4 Statistical Validation of $rep$ via Strength Predicates
Recall that since the semantics of $\texttt{Con}_{\texttt{spec}}$ are parameterized by the concept representation map $rep$, the validity of the semantics is dependent on the validity of $rep$. Our next step is to statistically validate the $rep$ implemented using CLIP model $g$ (as per Defn. 3) as well the one implemented via the affine map $r_{map}$ (as per Defn. 4) between the ResNet18 $f$ and $g$’s representations spaces.
Towards this end, we statistically validate our formulated strength predicates using the RIVAL10 test data ($D_{\text{test}}$) for the concept representation maps under consideration. In particular, for a given $rep$, we measure the rate (referred to as satisfaction probability), for all classes $c$, at which inputs in $D_{\text{test}}$ with ground-truth label $c$ satisfy the strength predicates formulated for $c$. Our intuition is that, given an ideal concept representation map, the strength predicates for a class $c$ should always hold, i.e., with probability 1.0 (unless the formulated predicates are nonsensical, for instance, $hairy\texttt{>}metallic$ for class truck). Thus, consistently low satisfaction probabilities indicate a low-quality $rep$. Note that this procedure neither requires access to an ideal concept representation map nor data with manually annotated concept labels if the strength predicates are given.
@@ Results.
The results for $rep$ implemented only using CLIP model $g$ (as in Defn. 3) are shown in Fig. 5(a) and 5(b).Concepts on the Y-axis represent $cor_{r}$ and on the X-axis represent $con_{ir}$ in the evaluated strength predicates. We see that for class truck, except for strength predicates involving concept text, the satisfaction probabilities are consistently high. Similar results are observed for class car as well as other classes. This provides strong evidence for our hypothesis that a concept representation implemented using CLIP is of high-quality.
The results for $rep$ implemented using the ResNet18 model $f$ and CLIP model $g$ via the affine map (as per Thm. 4.1) are summarized in Fig. 6(a) and 6(b). The trends are similar to those observed for $rep$ implemented only using $g$. In addition to suggesting that the $rep$ based on the affine map $r_{map}$ is also of high-quality, these results are evidence in favor of the assumption that the representation space of $f_{enc}$ is well aligned with the representation space of $g_{enc}^{img}$ via our learned affine map (also suggested by the low MSE and high R2 of $r_{map}$). However, note that the representation spaces are not faithfully aligned (for all inputs as per Defn. 5), as indicated by the counterexamples in Fig. 7.
@@ 5.5 Verification
Next, we use the elicited predicates to form specifications that we attempt to formally verify. In particular, we performed two sets of experiments—first, we conduct formal verification of the CLIP model itself, and then, we formally verify the ResNet18 model. In both cases, we can reduce the verification to solving of linear constraints (despite the fact that $cos$ involves non-linear constraints). For space reasons, we report only the ResNet18 results here, while the CLIP results are delegated to Appendix 0.D.
We restrict our investigation only to statistically significant concept predicates, i.e., those that hold on the train set with over $95\%$ probability, since formal proofs are more likely in such cases. Moreover, in order to decide if a model satisfies a specification, we need to define an input scope $B\subseteq X$ (see Defn. 1).
Ideally, we would like to show that a specification holds for all inputs $v\in X$ but we find that such a requirement is too strong. Majority of the inputs in set $X$ do not correspond to any meaningful image (i.e., are out-of-distribution), and attempting to verify model behavior at such inputs typically leads to verification failure without discovering any useful counter-examples.
An effective compromise would be to verify the models on inputs that are in-distribution, i.e., in the support set of the distribution over $X$ (denoted as $\textit{inDist}(X)$) that characterizes the input data. While a sound and complete encoding of the inDist predicate (i.e., an encoding that evaluates to True if and only if a sample is in-distribution) is not feasible, we describe simple approximations of the same that can be useful in practice. We also leverage Thm. 4.1 to conduct proofs in the representation space instead of the input space.
@@
For this experiment, we verify a ResNet18 vision model $f$ while using CLIP as a VLM $g$ to implement the concept representation map $rep$ as per Defn. 4. We first identify interesting regions in the embedding space of the network, and then check if the statistically significant concept predicates hold in these regions.
@@
We want to formally check if the vision model $f$ satisfies a specification $e$ with respect to a concept representation map $rep$ (as defined in Defn. 4) and an input scope $B$ (i.e., $(f,rep,B)\models e$) which, using Thm. 4.1, can be rephrased as $(f_{head},\widehat{rep},\overline{B})\models e$).
@@
To define the concept representation map $\widehat{rep}$ as per Defn. 4, we first extract embeddings corresponding to concepts in the embedding space of CLIP (denoted as $q^{con_{i}}$ for concept $con_{i}$) using the CLIP text encoder. As explained in Section 4, we also learn an affine map $r_{map}$ such that $z=r_{map}(w)$ where $w$ is a vector in the vision model $f$’s embedding space and $z$ denotes a vector in the VLM model $g$’s embedding space. For simplicity, we assume that the length of both vectors is $p$ which holds in this case.
@@ 6 Related Work
Formal Analysis of DNNs. The formal analysis of DNNs, especially in safety-critical applications, has been a subject of growing interest. Works like [19, 21, 12] have explored methods to verify the safety and correctness of DNNs, but they often grapple with the complexity and scale of these networks. Our approach differs by translating the problem into the realm of natural language, thus potentially bypassing some of the direct complexities involved in analyzing the networks themselves. Further, existing DNN verification methods are restricted to simple robustness properties while our approach enables semantic specification and verification of DNN models.
@@ 7 Conclusion
While in this paper we mainly focused on checking specifications as formal explanations for DNN decisions, in the future we plan to explore other uses of $\texttt{Con}_{\texttt{spec}}$ specifications, e.g., as run-time checks to detect mis-classifications or adversarial attacks. An important open challenge is formally defining $B$ to only include in-distribution inputs and avoid noise. Another direction we intend to explore is to use metrics other than cosine similarity for comparing embeddings. We also plan to experiment with more multimodal models and to assess the effectiveness of our techniques in safety-critical applications, which have clear definition of concepts.
Appendix 0.A Faithful Alignment of Representation Spaces
Defn. 5 defines the notion of a faithful representation space alignment. Intuitively, the alignment is faithful if no information is lost when using $r_{map}$ to map between the two representation spaces.
Definition 5 (Faithful alignment of representation spaces)
$\forall x\in X.~{}r_{map}(f_{enc}(x))=g_{enc}^{img}(x)$
Appendix 0.B Proofs
Lem. 1,  2, and Thm. 4.1 are all proven for a vision model $f:X\rightarrow Y$ that can be decomposed into $f_{enc}:X\rightarrow Z$ and $f_{head}:Z\rightarrow Y$ where $Z:=\mathbb{R}^{d^{\prime}}$, a $\texttt{Con}_{\texttt{spec}}$ specification $e$, a linear concept representation map $rep$ (as defined in Defn. 2 or 4), and an input scope $B\subseteq X$. They use the notation $\widehat{rep}$ for the function obtained from $rep$ by replacing $f_{enc}$ with the identity function, $f_{enc}(B):=\{f_{enc}(x)\mid x\in B\}$, $f_{enc}^{-1}(v^{\prime}):=\{v\mid f_{enc}(v)=v^{\prime}\}$ for the preimage under $f_{enc}$ of a point $v^{\prime}$, and IH for induction hypothesis.
A general observation that holds given the definitions of $rep$ (as per Defn. 2 or 4) and $\widehat{rep}$ (all occurrences of $f_{enc}$ in $rep$ replaced by the identity function) is that,
@@ Appendix 0.B Proofs
Theorem 0.B.1
Given a VLM $g$ with encoders $g_{enc}^{img}:X\rightarrow Z$ and $g_{enc}^{txt}:T\rightarrow Z$ and head $g_{head}:Z\rightarrow Y$ where $Z:=\mathbb{R}^{d^{\prime}}$, a $\texttt{Con}_{\texttt{spec}}$ specification $e$, a linear concept representation map $rep$ (as defined in Defn. 3), and an input scope $B\subseteq X$,
@@
Appendix 0.C Captions Used for Computing CLIP Text Embeddings
@@
Appendix 0.D Verification of CLIP
@@
We want to formally check if the CLIP model $(g_{head}\circ g_{enc}^{img})$ satisfies a specification $e$ with respect to a concept representation map $rep$ and an input scope $B$ (i.e., $(g_{head}\circ g_{enc}^{img},rep,B)\models e$) which, using Thm. 0.B.1, can be rephrased as $(g_{head},\widehat{rep},\overline{B})\models e$).
We use the zero-shot classification method as described in Section 2 to define $g_{head}$ for determining the most probable class. We compute the mean text embedding $q^{c_{j}}$ for each class $c_{j}$ (using the CLIP text encoder $g_{enc}^{txt}$ and the set of captions listed in Appendix 0.C) and calculate $cos$ similarity between an input image embedding and these classes’ text embeddings. The closest class embedding is the predicted class. We similarly compute the embedding corresponding to each concept ${con_{i}}$, denoted $q^{con_{i}}$ here and as $\overline{con_{i}}$ in Defn. 3, in CLIP’s embedding space. These embeddings represent the direction corresponding to the concepts and are used to define $\widehat{rep}$.

### v3

@@ 5.4 Statistical Validation of $rep$ via Strength Predicates
Recall that since the semantics of $\texttt{Con}_{\texttt{spec}}$ are parameterized by the concept representation map $rep$, the validity of the semantics is dependent on the validity of $rep$. Our next step is to statistically validate the $rep$ implemented using CLIP model $g$ (as per Defn. 2) as well the one implemented via the affine map $r_{map}$ (as per Defn. 3) between the ResNet18 $f$ and $g$’s representations spaces.
Towards this end, we statistically validate our formulated strength predicates using the RIVAL10 test data ($D_{\text{test}}$) for the concept representation maps under consideration. In particular, for a given $rep$, we measure the rate (referred to as satisfaction probability), for all classes $c$, at which inputs in $D_{\text{test}}$ with ground-truth label $c$ satisfy the strength predicates formulated for $c$. Our intuition is that, given an ideal concept representation map, the strength predicates for a class $c$ should always hold, i.e., with probability 1.0 (unless the formulated predicates are nonsensical, for instance, $hairy\texttt{>}metallic$ for class truck, represented by yellow-colored squares in Fig. 5 and 6). Thus, consistently low satisfaction probabilities indicate a low-quality $rep$. Note that this procedure neither requires access to an ideal concept representation map nor data with manually annotated concept labels if the strength predicates are given.
@@ Results.
The results for $rep$ implemented only using CLIP model $g$ (as in Defn. 2) are shown in Fig. 4(a) and 4(b). Concepts on the Y-axis represent $cor_{r}$ and on the X-axis represent $con_{ir}$ in the evaluated strength predicates. We see that for class truck, except for strength predicates involving concept text, the satisfaction probabilities are consistently high. Similar results are observed for class car as well as other classes. This provides strong evidence for our hypothesis that a concept representation implemented using CLIP is of high-quality.
[…]
@@ 5.5 Verification
Next, we use the elicited predicates to form specifications that we attempt to formally verify. In particular, we performed two sets of experiments—first, we conduct formal verification of the CLIP model itself, and then, we formally verify the ResNet18 model. In both cases, we can reduce the verification to solving of linear constraints (despite the fact that $cos$ involves non-linear constraints). Note that the reduction to linear constraints relies on the assumption that the head of the model is a linear layer. While this assumption holds true for the models we consider, in general the head may have non-linear layers in which case we would need a more powerful solver to address the non-linear constraints. For space reasons, we report only the ResNet18 results here, while the CLIP results are delegated to Appendix 0.C.
We restrict our investigation only to statistically significant strength predicates, i.e., those that hold on the train set with over $95\%$ probability, since they are more likely to be semantically meaningful and are more likely to hold. Moreover, in order to decide if a model satisfies a specification, we need to define an input scope $B\subseteq X$ (see Defn. 1).
We could try to show that a specification holds for all inputs $v\in X$ but we find that such a requirement is too strong. Majority of the inputs in set $X$ do not correspond to any meaningful image (i.e., are out-of-distribution), and attempting to verify model behavior at such inputs typically leads to verification failure without discovering any useful counter-examples.
Ideally, we would like to verify the models on inputs that are in-distribution, i.e., in the support set of the distribution over $X$ that characterizes the input data. Unfortunately, this is an unsolved challenge in the research community, as it is not known how to mathematically represent in-distribution data in a sound and complete manner. We describe here approximations of the same that can be useful in practice. We also leverage Thm. 4.1 to conduct proofs in the representation space instead of the input space.
@@
For this experiment, we verify a ResNet18 vision model $f$ while using CLIP as a VLM $g$ to implement the concept representation map $rep$ as per Defn. 3. We first identify interesting regions in the embedding space of the network, and then check if the statistically significant concept predicates hold in these regions.
@@
We want to formally check if the vision model $f$ satisfies a specification $e$ with respect to a concept representation map $rep$ (as defined in Defn. 3) and an input scope $B$ (i.e., $(f,rep,B)\models e$) which, using Thm. 4.1, can be rephrased as $(f_{head},\widehat{rep},\overline{B})\models e$).
@@
To define the concept representation map $\widehat{rep}$, we first extract embeddings corresponding to concepts in the embedding space of CLIP (denoted as $q^{con_{i}}$ for concept $con_{i}$) using the CLIP text encoder. As explained in Section 4, we also learn an affine map $r_{map}$ such that $z=r_{map}(w)$ where $w$ is a vector in the vision model $f$’s embedding space and $z$ denotes a vector in the VLM model $g$’s embedding space. For simplicity, we assume that the length of both vectors is $p$ which holds in this case.
@@ 6 Related Work
Formal Analysis of DNNs. The formal analysis of DNNs, especially in safety-critical applications, has been a subject of growing interest. Works like [19, 21, 12] have explored methods to verify the safety and correctness of DNNs, but they often grapple with the complexity and scale of these networks. Our approach differs by translating the problem into the realm of natural language, thus potentially bypassing some of the direct complexities involved in analyzing the networks themselves. Further, existing DNN verification methods are restricted to simple robustness properties expressed in terms of input pixels while our approach enables semantic specification and verification of DNN models.
@@ 7 Conclusion
While in this paper we mainly focused on checking specifications as formal explanations for DNN decisions, in the future we plan to explore other uses of $\texttt{Con}_{\texttt{spec}}$ specifications, e.g., as run-time checks to detect mis-classifications or adversarial attacks. An important open challenge is formally defining $B$ to only include in-distribution inputs and avoid noise. One idea is to check the specifications at run-time; $\texttt{Con}_{\texttt{spec}}$ may be particularly useful for run-time monitoring of DNNs in safety-critical settings. Another direction we intend to explore is to use metrics other than cosine similarity for comparing embeddings. We also plan to experiment with more multimodal models and to assess the effectiveness of our techniques in safety-critical applications, which have clear definition of concepts.
Acknowledgements
This work was supported in part by the United States Air Force and DARPA under Contract No.FA8750-23-C-0519, and the U.S. Army Research Laboratory Cooperative Research Agreement W911NF-17-2-0196. Any opinions, findings and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the Department of Defense or the United States Government.
Appendix 0.A Proofs
Lem. 1,  2, and Thm. 4.1 are all proven for a vision model $f:X\rightarrow Y$ that can be decomposed into $f_{enc}:X\rightarrow Z$ and $f_{head}:Z\rightarrow Y$ where $Z:=\mathbb{R}^{d^{\prime}}$, a $\texttt{Con}_{\texttt{spec}}$ specification $e$, a linear concept representation map $rep$ (as defined in Defn. 3), and an input scope $B\subseteq X$. They use the notation $\widehat{rep}$ for the function obtained from $rep$ by replacing $f_{enc}$ with the identity function, $f_{enc}(B):=\{f_{enc}(x)\mid x\in B\}$, $f_{enc}^{-1}(v^{\prime}):=\{v\mid f_{enc}(v)=v^{\prime}\}$ for the preimage under $f_{enc}$ of a point $v^{\prime}$, and IH for induction hypothesis.
A general observation that holds given the definitions of $rep$ (as per Defn. 3) and $\widehat{rep}$ (all occurrences of $f_{enc}$ in $rep$ replaced by the identity function) is that,
@@ Appendix 0.A Proofs
Theorem 0.A.1
Given a VLM $g$ with encoders $g_{enc}^{img}:X\rightarrow Z$ and $g_{enc}^{txt}:T\rightarrow Z$ and head $g_{head}:Z\rightarrow Y$ where $Z:=\mathbb{R}^{d^{\prime}}$, a $\texttt{Con}_{\texttt{spec}}$ specification $e$, a linear concept representation map $rep$ (as defined in Defn. 2), and an input scope $B\subseteq X$,
@@
Appendix 0.B Captions Used for Computing CLIP Text Embeddings
@@
Appendix 0.C Verification of CLIP
@@
We want to formally check if the CLIP model $(g_{head}\circ g_{enc}^{img})$ satisfies a specification $e$ with respect to a concept representation map $rep$ and an input scope $B$ (i.e., $(g_{head}\circ g_{enc}^{img},rep,B)\models e$) which, using Thm. 0.A.1, can be rephrased as $(g_{head},\widehat{rep},\overline{B})\models e$).
We use the zero-shot classification method as described in Section 2 to define $g_{head}$ for determining the most probable class. We compute the mean text embedding $q^{c_{j}}$ for each class $c_{j}$ (using the CLIP text encoder $g_{enc}^{txt}$ and the set of captions listed in Appendix 0.B) and calculate $cos$ similarity between an input image embedding and these classes’ text embeddings. The closest class embedding is the predicted class. We similarly compute the embedding corresponding to each concept ${con_{i}}$, denoted $q^{con_{i}}$ here and as $\overline{con_{i}}$ in Defn. 2, in CLIP’s embedding space. These embeddings represent the direction corresponding to the concepts and are used to define $\widehat{rep}$.